# Cutting through buggy adversarial example defenses: fixing 1 line of code breaks Sabre

### v2

@@ Abstract
Sabre is a defense to adversarial examples that was accepted at IEEE S&P 2024. We first reveal significant flaws in the evaluation that point to clear signs of gradient masking. We then show the cause of this gradient masking: a bug in the original evaluation code. By fixing a single line of code in the original repository, we reduce Sabre’s robust accuracy to 0%. In response to this, the authors modify the defense and introduce a new defense component not described in the original paper. But this fix contains a second bug; modifying one more line of code reduces robust accuracy to below baseline levels. After we released the first version of our paper online, the authors introduced another change to the defense; by commenting out one line of code during attack we reduce the robust accuracy to 0% again.
@@ 1 Introduction
In this short paper we show that by changing a single line of code to fix a bug in the published evaluation code, the attack success rate increases to 100%. The authors responded to our attack by modifying the published defense code to introduce a new defense component that was not present in the original paper, thereby changing the defense to a different defense than was initially proposed. Unfortunately, this modification contains a second bug, and by modifying a second line of code we again completely break the defense. We then released the first version of our paper online. After this initial release of this paper, the authors updated the code again with another fix; but this contained a third bug, and by commenting out a single line code when running the attack (and putting it back at test time), we reduced defense accuracy to 0%.

### v3

@@ Abstract
Sabre is a defense to adversarial examples that was accepted at IEEE S&P 2024. We first reveal significant flaws in the evaluation that point to clear signs of gradient masking. We then show the cause of this gradient masking: a bug in the original evaluation code. By fixing a single line of code in the original repository, we reduce Sabre’s robust accuracy to 0%. In response to this, the authors modify the defense and introduce a new defense component not described in the original paper. But this fix contains a second bug; modifying one more line of code reduces robust accuracy to below baseline levels. After we released the first version of our paper online, the authors introduced another change to the defense; by commenting out one line of code during attack we reduce the robust accuracy to 0% again. Then, in response to this attack, the authors released another change to the code, reverting many changes and introducing new defense components—again; we break this version of the code by adding a single minus sign to one of the attack hyperparmaeters.
@@ 1 Introduction
In this short paper we show that SABRE’s impossible levels of accuracy can be explained by a bug in the original evaluation; fixing the bug by changing a single line of code increases the attack success rate increases to 100%. In response, in a process that will repeat three more times, the authors make a change to the code and introduce defense components never described in the original paper. We show that each of these three changes contains bugs that invalidate the results we break: the first by inserting a single function call; the second by commenting out a single line of code; and the third by negating one attack hyperparameter.
@@ 6.4 Our third break
6.5 Our fourth break
On June 4th the authors released a fourth version of their code in response to our prior attack. The purpose of this change, taken from the git commit title, is “refactoring with no discretization”. That is, it it completely removes all discretization steps that were introduced in the second and third revisions of the code. [^1] It also makes many other changes, again without explanation for how these changes relate to content described in the paper. Because these other changes do not impact our results we omit any further discussion. When asked if this was intentional, the authors stated that it was; and they assert this latest version of the code accurately reflects the experiments performed in the original paper. At the same time, in prior communication the authors asserted on multiple occasions that the discretization was a necessary component of their defense that was unintentionally omitted from the first code release.
Both of these can not be true simultaneously. The authors have now asserted that the code used in the experiments both did—and did not—include discretization. In this section we briefly study whether or not this version of the code could have been the code used for the original experiments in the paper. Then, we show that (regardless of whether or not this code was original) this implementation is also not robust to attack.
Unreproducible.
We find that this specific version of the code can not have been what was used to perform the original experiments: without any modifications, when we run the provided defense evaluation with the reference PGD attack on a CIFAR-10 model trained by the authors we find that it increases the defense accuracy to 100%. This is despite the original paper stating that running PGD on a CIFAR-10 trained model should yield 93% accuracy. (We are unable to run the MNIST model because the authors’ changes have since broken this model and its clean accuracy is now 10%.)
Our attack.
Nevertheless, we now break this version of the defense awaiting any future revisions. Our attack here is entirely trivial: whereas previously we have §6.1 removed a function call, §6.2 inserted a function call, and §6.3 commented out a line of code, in this section we now insert a single minus sign.
Specifically, we make the following observation: if performing PGD stepping in the direction of the gradient increases accuracy when under attack—why shouldn’t we just step in the opposite direction of the gradient to decrease accuracy when under attack? To do this, all we need to do is negate the value of $\alpha$: the attack’s step size hyperparameter.
1diff --git a/notebooks/cifar10.ipynb b/notebooks/cifar10.ipynb
2index 543aa90..f1e5553 100644
3--- a/notebooks/cifar10.ipynb
4$+$ b/notebooks/cifar10.ipynb
5@@ -389,7 +389,7 @@
6 "metadata": {},
7 "outputs": [],
8 "source": [
[…]
We find this attack completely breaks the latest version of the code: negating the step direction reduces defense accuracy to 4%, well below the random guessing baseline. An explanation of this phenomenon is beyond the scope of this paper. It could be because the model was retrained (again) in a different way and this retraining caused gradients to point backwards; or it could be because the reconstruction model now behaves differently than it did previously.
Given the simplicity of this attack, we aim to confirm it is effective by providing the authors with a list of adversarial examples we generated by negating the step size. Despite several attempts, the authors have been unwilling to verify the efficacy of these adversarial examples.